3.4 A storage tank is 5 m in diameter, 4 m high and the density of the fluid it

contains is 795 kg/m3. Determine the required wall thickness if the yield

stress of the material is 325 MPa, the factor of safety is 10 and the efficiency of

the longitudinal joint is 80 %.


1
Expert's answer
2021-03-19T06:48:19-0400

tm = Fs Pmdi /m

tm = Minimum wall thickness (m)

Fs = Safety factor

Pm = Internal pressure (N/m2)

di = Inside diameter (m)

δm = Metal yield stress (N/m2)

Pm = "\rho" g h = (795 kg/m3)(9.8 m/s2)(4 m) = 31 164 N/m2

tm = 10 (31 164 N/m2)(5 m)/ 2(325 000 000 Pa)0.8 = 0.0030 m = 3.0 mm